A fishpond is 12 meters long by 8 meters wide. The owner wants to double the are of the pond by increasing the length and width by the same amount. By how much should the length and width be increased?
Answer by macston(5194)

L=original length=12 meters; W=original width=8 meters; X=increase in size
Original Area=LW=(12m)(8m)=96m^2
New area =2(Original Area)=2(96m^2)=192m^2
.
(L+X)(W+X)=192m^2
(12m+X)(8m+X)=192M^2
X^2+20X+96=192
x^2+20X-96=0
(x+24)(x-4)=0
x+24=0 or x-4=0
x=-24m or x=4m
Since x=-24m doesn't make sense,
x=4 meters is the increase in lenght and width.
.
CHECK:
(L+X)(W+X)=2LW
(12m+4m)(8m+4m)=2(12m)(8m)
(16m)(12m)=2(96m^2)
192m^2=192m^2